OpenStudy (anonymous):

Logan is buying equipment for the football team and the baseball team. He spends $40 on football uniforms, and he plans on purchasing some footballs. Logan uses this equation to show the total cost of football equipment, y, based on the number of footballs purchased, x. y = 40 + 10x Logan also spends $200 on baseball uniforms, and he plans on purchasing some baseballs. Logan uses this equation to show the total cost of baseball equipment, y, based on the number of baseballs purchased, x. y = 200 + 2x Logan wants to spend the same amount of money on the football team and the base

OpenStudy (texaschic101):

40 + 10x = 200 + 2x 10x - 2x = 200 - 40 8x = 160 x = 20 if he buys 20 of each, he will have spent $ 240 on each.